OpenStudy (anonymous):

Find the slope of the line whose equation is 8y = 2x + 4. 4 1/2 correct? 1/4

OpenStudy (a1234):

8y = 2x + 4 Divide each side by 8: 8y/8 = 2x/8 + 4/8 y = 1/4x + 1/2 No, the slope is 1/4.

OpenStudy (a1234):

1/2 is the y-intercept.